For a normal eye the far point is infinity and the near point is at 25 cm in front of the eye. Cornea provides converging power of about 40D and the least converging power of eye lens behind the cornea is about 20D. Estimate the range of accomodation of normal eye.
1
Expert's answer
2013-02-15T04:40:58-0500
The range of accommodation isdefined as the distance between the farthest point at which an object can be seen clearly with accommodation fully relaxed, and the nearest distance at which an object can be observed with full accommodation.
Least converging power of cornea and the lens is 40 D + 20 D = 60 D; the corresponding focus is 1 / (60 D) = 1.66 cm.
The converging power of the cornea itself is 40 D, the corresponding focus is 1 / (40 D) = 2.5 cm.

The difference is 2.5 cm - 1.66 cm = 0.84 cm.

Answer: 0.84 cm